OpenStudy (dtan5457):

simplify the cube root of 32x^6 y^4

OpenStudy (dtan5457):

are you rewriting my question?

OpenStudy (dtan5457):

you wrote my question wrong

OpenStudy (dtan5457):

its 32x but its ok go on

OpenStudy (anonymous):

\[\sqrt[3]{32 x^6y^4}=\left( 32 x^6 y^4 \right)^{\frac{ 1 }{ 3 }}=\left( 2^3*4\left( x^2 \right) ^3y^3y \right)^{\frac{ 1 }{ 3 }}\] =?

OpenStudy (dtan5457):

Already figured this out a more easier way. Thanks for the help though.